# Reaction Prediction via Interaction Modeling of Symmetric Difference Shingle Sets

### v2

@@ Appendix B Datasets statistics
We use two datasets for pre-training and seven datasets for downstream evaluation, as summarized in Table 3. We remove duplicate records and invalid reactions for pre-training by RDKit [rdkit]. To assess the generalizability of our approach, we consider both random and out-of-sample splits. In the out-of-sample split, the test set contains reactions involving molecules that do not appear in the training set. This setup mimics real-world scenarios, where unseen molecular structures are encountered during reaction property prediction. To strike a balance between accuracy in spatial coordinates and computational efficiency, we utilize the ETKDG algorithm [ETKDG] to generate up to 100 conformers and sample one for each molecule.
@@
  • The United States Patent and Trademark Office (USPTO) dataset [USPTO] was collected from 1976 to September 2016, containing over 1.8 million chemical reactions stored in the form of SMILES arbitrary target specification (SMARTS).
@@
  • The Chemical Journals with High Impact Factor (CJHIF) dataset [CJHIF] included over 3.2 million chemical reactions in the form of SMARTS extracted from chemistry journals by Chemical.AI.
@@
  • The Buchwald-Hartwig (BH) dataset [rea_BH] contained 3,955 reactions from high-throughput experiments (HTEs) with 1,536-well plates on the class of Pd-catalyzed Buchwald-Hartwig C-N cross-coupling reactions. We choose the pyridyl reactants as the pivot to construct the out-of-sample split condition, where training includes nine pyridyl reactants and testing uses three non-pyridyl reactants, as shown in Figure 8(a).
@@
  • The Suzuki-Miyaura (SM) dataset [rea_SM] was constructed from high-throughput experiments on the class of Suzuki-Miyaura cross-coupling reactions, resulting in measured yields for a total of 5,760 reactions. We choose a set of ligands as the pivot to construct the out-of-sample split condition, where nine ligands (including “None”) are used for training and three for testing, as shown in Figure 8(b).
@@
  • The Ni-catalyzed C-O bond activation (NiCOlit) dataset [rea_NiCOlit] was extracted from organic reaction publications to form C-C and C-N bonds, containing 1,406 reactions. We choose the OPiv substrates as the pivot to construct the out-of-sample split condition, where 247 OPiv substrates are used for training and 42 non-OPiv substrates for testing, as shown in Figure 8(c).
@@
  • The real-world electronic laboratory notebook (ELN) dataset [rea_gnn_desc] was created for Buchwald-Hartwig reactions from electronic laboratory notebooks, including 750 reactions. The structural diversity of the real-world ELN dataset is much higher than that of the HTE datasets. The random split also simulates an out-of-sample scenario.
@@
  • The asymmetric N,S-acetal formation using CPA catalysts (N,S-acetal) dataset [rea_NS] included combinatorial variations of CPA catalysts, N-acyl imines, and thiols, resulting in a total of 1,075 reactions. We choose a set of catalysts (test-cat) as the pivot to construct the out-of-sample split condition, including 24 catalysts for training and 19 for testing, as shown in Figure 8(d).
@@
  • The nanomole-scale reactivity evaluation of C-heteroatom-coupling reactions (C-heteroatom) dataset [rea_heteroatom] was performed by an automated high-throughput screening on a nanomole scale, yielding 1,536 reactions. We choose a set of catalysts (test-cat) as the pivot to construct the out-of-sample split condition, where the split uses 12 catalysts for training and 4 for testing, as shown in Figure 8(e).
@@
  • USPTO_TPL dataset [mapping] labels were generated by extracting the 1,000 most common templates from the USPTO dataset, containing 445,115 reactions.
@@
USPTO [USPTO] & CJHIF [CJHIF] 3,728,503 stratified 3,542,077 186,426
BH [rea_BH] 3,955 random 2,768 1,187
@@
ELN [rea_gnn_desc] 750 random 525 225
N,S-acetal [rea_NS] 1,075 random 600 475
@@ Appendix D Pre-training settings
We introduce three pseudo-reaction-type classification tasks as pre-training objectives. The underlying intuition is that similar chemical structures tend to exhibit consistent semantic behavior across various reactions. Specifically, we employ the DRFP [rea_DRFP] with default parameters, which is a 1024-length one-hot descriptor as the reaction fingerprint. These reaction fingerprint sequences can be used for clustering, where shorter distances between fingerprints indicate a higher likelihood of belonging to the same cluster. We apply $K$ means clustering by scikit-learn [scikit-learn] with different values of $K$ (100, 1,000, 4,000, see Figure 9 concerning selection of $K$) to cluster reactions to obtain clusters with different granularities from coarse-grained to fine-grained. Subsequently, we use three classification heads with output sizes of 100, 1,000, and 4,000, respectively, to predict the pseudo labels associated with each cluster. This approach leverages the structural consistency of chemical reactions, enabling the model to learn more robust and transferable representations.
@@ Appendix E Comparison with standard cross-attention architecture
We implement a cross-attention baseline that explicitly models interactions between reactants and products. In this architecture, product embeddings serve as queries, while reactant embeddings are used as keys and values within a standard multi-head cross-attention module [transformer]. The model consists of 4 cross-attention layers, 64 attention heads, and a hidden dimension of 512, and it shares the same molecular encoders as ReaDISH.
@@ Appendix F Implementation details
Table 5 and 6 present hyper-parameters used in ReaDISH during pre-training and fine-tuning, respectively. We use Pytorch [pytorch] with the Adam [adam] optimizer and the cosine learning rate decay strategy for training. All experiments are executed on 4 NVIDIA RTX3090 GPUs. Pre-training the model takes about 8 hours, whereas fine-tuning on downstream datasets requires up to 1 hour. Generating shingles for the pre-training dataset takes around 7 hours using 50 CPU threads, while processing a downstream dataset completes in under 4 minutes. ReaDISH contains about 16.3M trainable parameters.

### v3

@@ Appendix B Datasets statistics
We use two datasets for pre-training and seven datasets for downstream evaluation, as summarized in Table 3. We remove duplicate records and invalid reactions for pre-training by RDKit [24]. To assess the generalizability of our approach, we consider both random and out-of-sample splits. In the out-of-sample split, the test set contains reactions involving molecules that do not appear in the training set. This setup mimics real-world scenarios, where unseen molecular structures are encountered during reaction property prediction. To strike a balance between accuracy in spatial coordinates and computational efficiency, we utilize the ETKDG algorithm [38] to generate up to 100 conformers and sample one for each molecule.
@@
  • The United States Patent and Trademark Office (USPTO) dataset [26] was collected from 1976 to September 2016, containing over 1.8 million chemical reactions stored in the form of SMILES arbitrary target specification (SMARTS).
@@
  • The Chemical Journals with High Impact Factor (CJHIF) dataset [20] included over 3.2 million chemical reactions in the form of SMARTS extracted from chemistry journals by Chemical.AI.
@@
  • The Buchwald-Hartwig (BH) dataset [1] contained 3,955 reactions from high-throughput experiments (HTEs) with 1,536-well plates on the class of Pd-catalyzed Buchwald-Hartwig C-N cross-coupling reactions. We choose the pyridyl reactants as the pivot to construct the out-of-sample split condition, where training includes nine pyridyl reactants and testing uses three non-pyridyl reactants, as shown in Figure 8(a).
@@
  • The Suzuki-Miyaura (SM) dataset [33] was constructed from high-throughput experiments on the class of Suzuki-Miyaura cross-coupling reactions, resulting in measured yields for a total of 5,760 reactions. We choose a set of ligands as the pivot to construct the out-of-sample split condition, where nine ligands (including “None”) are used for training and three for testing, as shown in Figure 8(b).
@@
  • The Ni-catalyzed C-O bond activation (NiCOlit) dataset [43] was extracted from organic reaction publications to form C-C and C-N bonds, containing 1,406 reactions. We choose the OPiv substrates as the pivot to construct the out-of-sample split condition, where 247 OPiv substrates are used for training and 42 non-OPiv substrates for testing, as shown in Figure 8(c).
@@
  • The real-world electronic laboratory notebook (ELN) dataset [40] was created for Buchwald-Hartwig reactions from electronic laboratory notebooks, including 750 reactions. The structural diversity of the real-world ELN dataset is much higher than that of the HTE datasets. The random split also simulates an out-of-sample scenario.
@@
  • The asymmetric N,S-acetal formation using CPA catalysts (N,S-acetal) dataset [62] included combinatorial variations of CPA catalysts, N-acyl imines, and thiols, resulting in a total of 1,075 reactions. We choose a set of catalysts (test-cat) as the pivot to construct the out-of-sample split condition, including 24 catalysts for training and 19 for testing, as shown in Figure 8(d).
@@
  • The nanomole-scale reactivity evaluation of C-heteroatom-coupling reactions (C-heteroatom) dataset [42] was performed by an automated high-throughput screening on a nanomole scale, yielding 1,536 reactions. We choose a set of catalysts (test-cat) as the pivot to construct the out-of-sample split condition, where the split uses 12 catalysts for training and 4 for testing, as shown in Figure 8(e).
@@
  • USPTO_TPL dataset [45] labels were generated by extracting the 1,000 most common templates from the USPTO dataset, containing 445,115 reactions.
@@
USPTO [26] & CJHIF [20] 3,728,503 stratified 3,542,077 186,426
BH [1] 3,955 random 2,768 1,187
@@
ELN [40] 750 random 525 225
N,S-acetal [62] 1,075 random 600 475
@@ Appendix D Pre-training settings
We introduce three pseudo-reaction-type classification tasks as pre-training objectives. The underlying intuition is that similar chemical structures tend to exhibit consistent semantic behavior across various reactions. Specifically, we employ the DRFP [35] with default parameters, which is a 1024-length one-hot descriptor as the reaction fingerprint. These reaction fingerprint sequences can be used for clustering, where shorter distances between fingerprints indicate a higher likelihood of belonging to the same cluster. We apply $K$ means clustering by scikit-learn [32] with different values of $K$ (100, 1,000, 4,000, see Figure 9 concerning selection of $K$) to cluster reactions to obtain clusters with different granularities from coarse-grained to fine-grained. Subsequently, we use three classification heads with output sizes of 100, 1,000, and 4,000, respectively, to predict the pseudo labels associated with each cluster. This approach leverages the structural consistency of chemical reactions, enabling the model to learn more robust and transferable representations.
@@ Appendix E Comparison with standard cross-attention architecture
We implement a cross-attention baseline that explicitly models interactions between reactants and products. In this architecture, product embeddings serve as queries, while reactant embeddings are used as keys and values within a standard multi-head cross-attention module [56]. The model consists of 4 cross-attention layers, 64 attention heads, and a hidden dimension of 512, and it shares the same molecular encoders as ReaDISH.
@@ Appendix F Implementation details
Table 5 and 6 present hyper-parameters used in ReaDISH during pre-training and fine-tuning, respectively. We use Pytorch [31] with the Adam [22] optimizer and the cosine learning rate decay strategy for training. All experiments are executed on 4 NVIDIA RTX3090 GPUs. Pre-training the model takes about 8 hours, whereas fine-tuning on downstream datasets requires up to 1 hour. Generating shingles for the pre-training dataset takes around 7 hours using 50 CPU threads, while processing a downstream dataset completes in under 4 minutes. ReaDISH contains about 16.3M trainable parameters.